### v3

# AverageTime: Enhance Long-Term Time Series Forecasting with Simple Averaging

Abstract
Long-term time series forecasting focuses on leveraging historical data to predict future trends. The core challenge lies in effectively modeling dependencies both within sequences and channels. Convolutional Neural Networks and Linear models often excel in sequence modeling but frequently fall short in capturing complex channel dependencies. In contrast, Transformer-based models, with their attention mechanisms applied to both sequences and channels, have demonstrated strong predictive performance. Our research proposes a new approach for capturing sequence and channel dependencies: AverageTime, an exceptionally simple yet effective structure. By employing mixed channel embedding and averaging operations, AverageTime separately captures correlations for sequences and channels through channel mapping and result averaging. In addition, we integrate clustering methods to further accelerate the model’s training process. Experiments on real-world datasets demonstrate that AverageTime surpasses state-of-the-art models in predictive performance while maintaining efficiency comparable to lightweight linear models. This provides a new and effective framework for modeling long time series.
Machine Learning, ICML
1 Introduction
Long-term time series prediction involves forecasting future trends over extended periods based on historical changes. This approach is crucial in various fields such as weather forecasting (Ma et al., 2023), traffic prediction (Long et al., 2024), and power demand estimation (Baruque et al., 2019). The extended forecast horizon and the complex correlations across sequences and channels pose significant modeling challenges. Traditional methods often fall short of capturing sequence and inter-channel relationships. Consequently, the primary methodologies in this field have shifted to deep learning models.
The core issue in long-term time series analysis is extracting dependencies within sequences and channels, which are crucial for improving model performance and robustness in multi-channel prediction. Various methods have been developed to capture these correlations from time series data. Widely employed techniques include Transformers (Wu et al., 2021; Zhang & Yan, 2022; Zhou et al., 2022, 2021; Nie et al., 2023; Liu et al., 2024; Wang et al., 2024b), which apply attention mechanisms to capture correlations both within sequences and across channels; Convolutional Neural Networks (CNN) (Wang et al., 2022; Wu et al., 2023), which use 1D or multidimensional convolutions to capture these dependencies; and structures based on Multilayer Perceptrons(Zeng et al., 2023; Chen et al., 2023; Lin et al., 2024b, a; Wang et al., 2024a; Li et al., 2023), such as DLinear, which decompose sequences and apply multiple linear layers to predict future series. Among these, the iTransformer (Liu et al., 2024), which captures channel dependencies using channel embeddings and an attention structure, has achieved significant performance improvements on benchmark datasets, especially in datasets with a large number of channels. However, as the importance of channel correlations has been increasingly recognized (Zhao & Shen, 2024), many models have not been able to effectively capture these correlations, leading to a decline in model performance. This inadequacy becomes particularly evident in scenarios with a large number of channels, resulting in suboptimal performance.
In this work, we introduce a simple yet effective framework for long-term time series forecasting, AverageTime, which employs mixed-channel embedding, parameter-independent prediction, and prediction averaging before and after embedding to model time series data. Specifically, we first embed the channels of the input data with MLP and Transformer encoders to generate data with the same dimensionality as the original input. This process can be repeated several times to further extract features. Subsequently, we employ parameter-independent MLP layers for channel-wise prediction. Additionally, we introduce a clustering strategy to predict channels with high correlation together, thereby accelerating the process by shifting from channel-wise to group-wise predictions. The prediction results from the embedded channel vectors are first averaged and directly added to the predictions from the original data.
Experimental results indicate that AverageTime surpasses the performance of the current leading transformer architecture while maintaining efficiency comparable to lightweight linear models. The primary contributions of AverageTime can be summarized as follows.
  • We introduce an averaging framework for long time series to capture inter-channel correlations. Specifically, we average the predictions of the original sequence with those obtained after channel embedding, achieving favorable performance in long-term time series forecasting.
  • We enhance our approach by incorporating a channel clustering mechanism, where channels with high similarity share the same prediction structure to accelerate the model training process.
  • We develop a mixed-channel embedding layer that combines the distinct characteristics of Linear and Transformer layers to enhance the model’s ability to capture channel dependencies.
2 Problem Statement
In multivariate time series analysis, the objective is to solve the following problem: Given an input time series $X$ of dimensions $C\times T_{1}$, where $C$ is the number of channels and $T_{1}$ is the length of the time series. The future time period to be forecast is set to $T_{2}$. Therefore, a deep learning architecture is required to map the input time series $X$ to the predicted output $Y$ for $T_{2}$ future steps. This can be formalized as:
$Y=f(X),$ (1)
where $f$ is a mapping function defined by a deep learning model that predicts $Y$ based on historical observations $X$. The function $f$ can be implemented using various architectures, including CNN, MLP, transformers, and their variants. Several methods have been developed to capture time series characteristics and enhance predictions, primarily categorized as follows.
2.1 Transformer Models
The Transformer architecture, an advanced method for long-term time series forecasting, excels at capturing dependencies within sequences and channels through the attention mechanism. The attention mechanism allows the model to focus on different parts of the input sequence when generating the output, efficiently handling sequences of varying lengths. The attention mechanism(Vaswani et al., 2017) is defined as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (2)
here, $Q$, $K$, and $V$ are feature matrices from the input data, and $d_{k}$ is the dimensionality of the keys. Recent research on transformer architectures has introduced several effective methods. PatchTST (Nie et al., 2023) divides the sequence into segments and applies attention, capturing relationships between different parts. iTransformer (Liu et al., 2024), embeds the channels and applies the attention mechanism to capture the complex dependency among channels, achieving significant performance improvements, especially in scenarios with a large number of channels. Additionally, methods like Informer (Zhou et al., 2021) address complexities in time series applications of attention, with these models demonstrating promising results.
2.2 MLP Models
In recent years, MLPs have become essential for multivariate long-term time series forecasting. With their straightforward architecture and strong predictive capabilities, MLPs play a pivotal role in multivariate time series forecasting. TiDE(Das et al., 2023) introduces a residual architecture into its linear layers to enhance the encoding of historical time series data and relevant covariates. The encoded information is then decoded with future covariates taken into account to generate forecasts. The dense encoder and decoder used in TiDE are shown as follows:
$\hat{\phi}_{\tau}$ $=M_{\phi}\cdot\phi_{\tau}+c_{\phi},$ (3)
where $M_{\phi}$ represents weight matrices, while $c_{\phi}$ denotes bias terms associated with the input time series. Through this approach, TiDE enhances the accuracy and robustness of predictions by leveraging both past data patterns and anticipated future conditions, thereby addressing the complexities inherent in time series forecasting. SparseTSF(Lin et al., 2024a) employs a cross-period modeling approach, wherein input data are first downsampled before being mapped through linear layers. The mapped data is then upsampled to generate the predicted time series output. Utilizing a linear model with fewer than 1k parameters, this approach achieves commendable forecasting performance. Such methodologies allow linear models to tackle time series forecasting with a highly efficient architecture, promoting their application in multivariate long-term forecasting tasks.
2.3 CNN Models
CNN-based time series analysis primarily relies on convolution operations to capture dependencies across both sequences and channels. MICN(Wang et al., 2022) achieves this through a multi-scale branch structure designed to model different underlying patterns separately, each branch extracts local features via downsampled convolution and captures global correlations through isometric convolution. Moreover, MICN exhibits linear complexity with respect to sequence length, making it more efficient than other models when dealing with long sequences. TimesNet(Wu et al., 2023) introduces the TimesBlock, a module that transforms 1D time series into 2D tensors. This transformation enables the modeling of complex temporal variations through efficient 2D convolution kernels, thereby effectively capturing both intra-series periodicities and inter-period variations. Furthermore, this approach is applicable to a variety of tasks within time series analysis, facilitating broader applications in time series forecasting and anomaly detection.
3 Method
3.1 Overview
In this section, we first introduce the principal architecture of AverageTime, which is composed of MLPs for prediction, along with transformer encoders and MLPs for channel embedding. To capture the correlations among channels, we design a simple yet effective framework for long-term time series forecasting, as illustrated in Figure 1.
For the input time series, this approach aims to model channel dependencies using a combination of Transformer-based and MLP-based channel embedding layers, all the channel layers are shared across all time steps. As shown in the top of the figure, we first transpose the time series, apply multiple Transformer encoders and MLPs to capture the channel dependencies, the embedded time series is then transposed back, after which both the transposed time series and the original time series are processed using parameter-independent MLP layers. Finally, the outputs are averaged to obtain the final prediction. Additionally, we enhance the prediction process using a clustering method. Specifically, during prediction, a channel clustering approach is employed to ensure the channels with high correlations share the same MLP parameters. This transformation shifts channel-wise predictions to group-wise predictions, thereby improving model efficiency while maintaining predictive performance.
Figure 1: The overall architecture of the AverageTime model, which employs MLPs for predictions and incorporates mixed-channel embedding layer with shared parameters to capture inter-channel correlations.
Our model effectively leverages information both across channels and sequences through the use of a parameter-shared mixed-channel embedding layer and parameter-independent MLP prediction layers. It also incorporates the Revin regularization technique(Kim et al., 2021) to mitigate distributional shifts between sequences.
The core framework of our approach involves an averaging operation between the embedded results and the original predicted results. Averaging after prediction effectively decouples the channel and sequence capture processes, reducing the interaction between these two processes and thus improving the model’s performance.
3.2 LightAverageTime
To address the efficiency issues encountered when the model iterates over all channels in AverageTime, particularly in scenarios with a large number of channels, we propose LightAverageTime, which introduces an additional clustering block as a complementary solution. The clustering block calculates the correlations among all channels and assigns MLPs to channel groups rather than individual channels, as illustrated in Figure 2:
Figure 2: The model initially computes the spearman correlation coefficients between time series. Subsequently, it employs a community detection algorithm to cluster the channels, thereby obtaining time series that are enriched with grouping information.
For a long-term time series forecasting task, assuming we have an input time series $X^{C\cdot T}$, to calculate the correlation between different channels and ensure fairness, we first split the time series into ${X_{train}}^{C\cdot T_{train}}$, ${X_{test}}^{C\cdot T_{test}}$, ${X_{val}}^{C\cdot T_{val}}$, then we calculate spearman correlation on the training dataset, which is:
$r_{s}=1-\frac{6\sum_{i=1}^{n}{d_{i}}^{2}}{n{(n^{2}-1)}},$ (4)
where $r_{s}$ stands for the rank correlation, $d_{i}$ stands for the rank difference for different channels, $n$ stands for the number of channels in the train dataset. Through the application of spearman correlation, we can have an estimate of the similarity among channels $G^{C\cdot C}$, where $G^{i\cdot j}$ stands for the spearman correlation between channel $i$ and channel $j$, we then apply a method to filter the channels with low correlation with all other channels, in a manner similar to the behavior of an activation function, which is:
[…]
where threshold is a hyperparameter that we define to determine how a channel differs from all other channels. During the main experiment, the threshold is uniformly set to 0.8. After that, we use the Label Propagation Algorithm(Zhou et al., 2003) to group the channels. We first initialize a label for each channel $L_{i}$, where $i=1,2,3,\dots,C$. Then, we iterate over the labels, updating each label by taking the majority of the neighbor labels. Specifically, for the $t_{th}$ iteration, the label for the $i_{th}$ channel is:
$L_{i}^{(t)}=\arg\max_{L_{j}\in\{L_{1},L_{2},\dots,L_{n}\}}\left(\sum_{j\in\mathcal{N}(i)}\delta(L_{j},L_{i})\right),$ (6)
where $L_{i}^{(t)}$ is the label of the $i_{th}$ channel for the $t_{th}$ iteration, $\mathcal{N}(i)$ represents the set of neighbors, denoted by matrix $G$, and $\delta(L_{j},L_{i})$ is the Kronecker delta function, which is defined as:
[…]
Through the computation of channel correlations and the application of community detection algorithms for grouping, our model enriches the time series channels with group information. The LightAverageTime approach then utilizes this channel group information to assign shared MLP prediction layers to channels within the same group. As a result, the iteration process shifts from iterating over individual channels to iterating over groups of channels, thereby increasing the speed of model operation.
4 Experiments
The experimental section is organized as follows: we begin by introducing the benchmark datasets we utilized, which are widely used in long-term time series forecasting. We then present the baselines used for comparison. Next, we describe the experimental setup and provide the main experimental results for AverageTime and LightAverageTime. Following this, we discuss the impact of hyperparameters on LightAverageTime as well as the effect of lookback windows on model performance. Afterward, we conduct a complexity analysis to assess the computational efficiency of our model. Finally, we present ablation studies to investigate the impact of the average framework on the model’s performance. All experiments in this section were conducted on a single NVIDIA RTX 4090D GPU.
4.1 Datasets
In this section, we first introduce the datasets used in our method, which include eight classic datasets for long-term time series forecasting. Their basic attributes are shown in Table 1:
Table 1: Datasets for long-term time series forecasting.
Data ETTh1 ETTh2 ETTm1 ETTm2 Weather Solar ECL Traffic
Channels 7 7 7 7 21 137 321 862
Frequency 1 hour 1 hour 15 mins 15 mins 10 mins 10 mins 1 hour 1 hour
Steps 17,420 17,420 69,680 69,680 52,696 52,179 26,304 17,544
Normally, the ETT datasets have fewer channels, making them easier to predict even with simple parameter-independent MLP layers. In contrast, for datasets with more channels, the average framework tends to be more effective due to its ability to effectively capture channel dependencies.
4.2 Baselines
We select several representative models for comparison: Transformer-based models : TimeXer, iTransformer, PatchTST, Linear models : CycleNet, RLinear, and CNN models : TimesNet. Note that we use CycleNet with MLP layers for better overall performance.
4.3 Experiments Setup
The basic setup for our AverageTime and LightAverageTime models is as follows: the look-back window is uniformly set to 96, and the prediction lengths are 96, 192, 336, and 720. Each model is trained for 30 epochs, with early stopping patience set to 5 to ensure the model is fully trained. The loss function used is Mean Squared Error, consistent with mainstream model training methods. The learning rate, batch size, and dropout rate vary depending on the dataset and prediction length. Note that long-term time series forecasting is highly sensitive to hyperparameters, and performance can be further improved through careful tuning.
4.4 Main results
In this section, we compare the primary results of AverageTime and LightAverageTime against the baseline models mentioned in Section 4.2 for long-term time series forecasting. Besides our proposed methods, the experimental results for other approaches are sourced from iTransformer (Liu et al., 2024) and TimeXer (Wang et al., 2024b). Specifically, for the Solar dataset, we implemented the TimeXer and CycleNet methods on a 4090D GPU ourselves and adjusted the hyperparameters to achieve better results, as their original papers did not provide these results. The results are then averaged over different prediction lengths. We bold the best and underline the second-best evaluation result for each dataset. The detailed comparison is presented in Table 2.
Table 2: Performance metrics for AverageTime and LightAverageTime in Multivariate Time Series.
Datasets Avg.Time LAvg.Time TimeXer CycleNet iTrans. RLinear PatchTST TimesNet
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 $\underline{0.438}$ 0.431 0.437 0.431 0.437 0.437 0.457 0.441 0.454 0.448 0.446 $\underline{0.434}$ 0.469 0.455 0.458 0.450
ETTh2 $\underline{0.372}$ $\underline{0.398}$ 0.377 0.401 0.368 0.396 0.388 0.409 0.383 0.407 0.374 0.399 0.387 0.407 0.414 0.427
ETTm1 $\underline{0.381}$ 0.391 0.383 $\underline{0.393}$ 0.382 0.397 0.379 0.396 0.410 0.410 0.414 0.408 0.390 0.400 0.398 0.406
ETTm2 $\underline{0.274}$ $\underline{0.318}$ 0.275 0.319 $\underline{0.274}$ 0.322 0.266 0.314 0.288 0.332 0.286 0.327 0.281 0.336 0.291 0.333
Weather 0.242 $\underline{0.271}$ 0.241 0.269 0.241 $\underline{0.271}$ 0.243 0.271 0.258 0.278 0.272 0.291 0.259 0.281 0.259 0.287
Solar $\underline{0.229}$ 0.271 0.237 0.273 0.233 0.281 0.210 $\underline{0.263}$ 0.233 0.262 0.369 0.357 0.254 0.289 0.301 0.319
ECL 0.174 $\underline{0.270}$ 0.176 0.272 $\underline{0.171}$ $\underline{0.270}$ 0.168 0.259 0.178 $\underline{0.270}$ 0.219 0.298 0.216 0.304 0.193 0.295
Traffic 0.424 0.279 $\underline{0.426}$ $\underline{0.281}$ 0.466 0.286 0.472 0.301 0.428 0.282 0.627 0.378 0.474 0.304 0.620 0.336
Average 0.317 0.329 $\underline{0.319}$ $\underline{0.330}$ 0.322 0.332 0.323 0.332 0.329 0.336 0.376 0.361 0.342 0.355 0.367 0.357
AverageTime achieved the best model performance among all compared models, with a 1.55% reduction in MSE and a 0.90% reduction in MAE compared with the second-best model, TimeXer, demonstrating the excellent performance of our model. The lightweight version, LightAverageTime with a cluster block, also outperforms TimeXer. Notably, the performance of AverageTime and TimeXer is nearly identical across the first seven datasets, with identical MSE performance and a 0.88% improvement in MAE. However, the mixed-channel embedding layer significantly improves our model’s performance on the Traffic dataset, indicating that AverageTime is more robust than TimeXer.
For datasets with a smaller number of channels, linear models like RLinear often exhibit sufficient predictive capability, as evidenced by the model’s superior performance on the ETT scenario. However, linear models typically fail in datasets with more channels. Moreover, our average framework is more effective for these scenarios. Notably, through mixed-channel embedding within the averaging framework, our model matched or even surpassed the performance of state-of-the-art Transformer architectures on Weather, ECL, and Traffic datasets, and the overall performance significantly outperforms Linear models like RLinear. This indicates that the average framework is simple yet of great importance in long-term time series forecasting.
4.5 Detail for LightAverageTime
As described in section 3.2, LightAverageLinear is highly dependent on the hyperparameter we define to filter out channels with low correlation to all other channels. For a fair comparison, we set the threshold value to 0.8 in the main results. The performance of LightAverageLinear can vary significantly depending on the selected threshold. When the threshold value approaches 1, the model behaves more like a parameter-independent method, whereas when the threshold is close to 0, it functions as a parameter-dependent method. Thus, adjusting the threshold can significantly affect the model’s predictions. The efficiency and performance of the model for different threshold values are illustrated in Figure 3:
(a) Different thresohold on model performance and efficiency on weather dataset.
(b) Different threshold on model performance and efficiency on electricity dataset.
Figure 3: Different threshold on model performance and efficiency on different datasets.
Note that the performance and training cost do not follow a simple linear relationship. By slightly reducing the threshold from 1, the training cost significantly decreases, while the performance suffers only a slight loss. This enables the model to select an appropriate threshold to achieve a suitable trade-off between training cost and performance. In the main experiments, LightAverageTime refers to a model with a threshold of 0.8. In practice, the threshold can be adjusted freely based on different requirements. We also include TimeXer and iTransformer for comparison. We first determine their performance in relation to the threshold, then calculate the efficiency at the corresponding threshold that our model needs to achieve comparable performance.
To be more specific, for weather datasets, when the threshold is around 0.8, LightAverageTime shows comparable performance to TimeXer, with only 67.5% of TimeXer’s training time and about half of the memory usage. Additionally, it continuously outperforms iTransformer across all thresholds. For electricity datasets, when the threshold is around 0.6, our model achieves the same performance as iTransformer, with 89.7% of the training time and 65% of the memory usage. Although our model slightly underperforms TimeXer, its training time is only 10% of TimeXer’s. This shows the flexibility of LightAverageTime and its great performance.
4.6 Complexity Analysis
To compare the complexity-related characteristics of different models, we first standardized several fundamental settings to ensure a fair comparison. Specifically, we set the batch size to 128 and the embedding dimension to 256. All other hyperparameters were set according to the default settings from their original papers. We evaluated the efficiency of all models on ETTh1 and weather datasets and recorded the training time per epoch and GPU memory usage for each model separately. The results of the complexity analysis are presented in Figure 4.
(a) ETTh1 efficiency.
(b) Weather efficiency.
Figure 4: Comparison of Model Complexities: We compare the complexities of different models. AverageTime and LightAverageTime demonstrate efficiency comparable to that of linear models, significantly outperforming Transformer-based models with approximately half the memory usage and 70% of the training time compared to the second-best model, TimeXer.
Notably, LightAverageTime achieves the best model efficiency on ETTh1 dataset, and second-best efficiency on the weather dataset, which is remarkable considering its great performance. It is followed by AverageTime, which outperforms other Transformer-based methods in terms of efficiency. Our models demonstrate overall efficiency comparable to that of linear models. In the weather dataset experiment, LightAverageTime can be executed faster than AverageTime due to the cluster block we designed, which switches the channel-wise prediction layers to group-wise prediction layers. Additionally, LightAverageTime is not restricted to a single model. While we set the threshold to 0.8 to ensure a fair comparison, reducing the threshold can further significantly improve the model’s efficiency.
4.7 Lookback windows experiment
Models with different lookback windows show varying performance. Some models that perform well with a lookback window of 96 may underperform with other lookback windows. To validate that the performance of AverageTime can be extended to different lookback windows, we conduct experiments on four datasets—ETTh1, ETTm1, weather, and traffic, with channel sizes ranging from 7 to 862. Hyperparameters are kept consistent with those used for a lookback window of 96 to simplify the experiment. The average results for different prediction lengths across the datasets are shown in Figure 5:
Figure 5: The model’s performance varies with changes in the lookback window.
It can be noted that, compared to the model with a lookback window of 96, the performance of AverageTime is consistently improved with longer lookback windows, which demonstrates that the average framework is effective across different lookback windows. The best performance is typically achieved when the prediction length equals 336. Note that the hyperparameters are consistent with those used for the model with a lookback window of 96. Since the model’s performance can be influenced by hyperparameters, carefully adjusting them can further enhance the model’s performance.
4.8 Ablation Analysis
In this section, we conduct an ablation study to evaluate the effectiveness of the averaging framework. Specifically, in the comparison model, we remove all channel embedding layers, leaving only simple MLP layers for prediction while keeping the hyperparameters consistent. The results are presented in Table 3.
Table 3: Ablation study on average framework.
Datasets MLP +Average Improv.(%)
Metrics MSE MAE MSE MAE MSE MAE
ETTh1 0.440 0.433 0.438 0.431 0.45 0.46
ETTh2 0.390 0.409 0.372 0.398 4.62 2.69
ETTm1 0.386 0.395 0.381 0.391 1.30 1.01
ETTm2 0.276 0.320 0.274 0.318 0.72 0.63
Weather 0.247 0.274 0.242 0.271 2.07 1.09
Solar 0.303 0.308 0.229 0.271 24.42 12.01
ECL 0.195 0.284 0.174 0.270 10.77 4.93
Traffic 0.517 0.325 0.424 0.279 17.99 14.15
Average 0.344 0.344 0.317 0.329 6.98 4.36
It can be observed that the averaging framework significantly enhances the model’s performance, particularly in datasets with a large number of channels, such as Solar, ECL, and Traffic. However, its effectiveness diminishes in datasets with fewer channels, though it still provides some improvement. This suggests that the impact of the averaging operation is more pronounced in high-channel settings and plays a crucial role in performance enhancement.
5 Conclusion
This paper introduces AverageTime, a method that outperforms state-of-the-art Transformer architectures like TimeXer, by employing an MLP-based prediction layer and a mixed-channel embedding layer, while maintaining efficiency comparable to that of lightweight linear models. We also propose LightAverageTime, which introduces a cluster block to group highly correlated channels for joint prediction, accelerating model training. This approach significantly enhances the computational efficiency of the model while maintaining robust predictive performance. Future work could explore the underlying mechanisms of the averaging framework for long-term time series forecasting and develop more advanced approaches to enhance predictive performance.
Appendix A Hyperparameter settings
Here, we provide the detailed hyperparameter settings for AverageTime. Specifically, we list the four key hyperparameters used in our experiments on the RTX 4090D: the number of transformer layers, the number of MLP layers, the dropout rate, and whether the model is channel-independent or channel-dependent. The detailed information is shown in Table 4:
Table 4: Hyperparameters for AverageTime.
Hyperparameters Transformer layers MLP layers Dropout rate Channel id
96 0 0 0.4 1
ETTh1 192 0 1 0.6 1
336 0 2 0.8 1
720 0 2 0.8 1
96 2 1 0.85 1
ETTh2 192 3 0 0.85 1
336 2 1 0.85 1
720 2 1 0.85 1
96 0 3 0.7 1
ETTm1 192 0 3 0.3 1
336 0 3 0.3 1
720 0 3 0.3 1
96 0 3 0.5 1
ETTm2 192 0 3 0.5 1
336 0 3 0.5 1
720 0 3 0.5 1
96 0 1 0 1
Weather 192 0 1 0 1
336 0 1 0 1
720 0 1 0.7 1
96 1 3 0 1
Solar 192 1 3 0 1
336 1 3 0 1
720 1 3 0 1
96 2 3 0 1
ECL 192 2 3 0 1
336 2 3 0 1
720 3 3 0 1
96 3 0 0 0
Traffic 192 3 0 0 0
336 3 0 0 0
720 3 0 0 0
Since hyperparameters play a crucial role in model performance, it is important to make them accessible and provide guidance on how to adjust them effectively.
For the ETTh2, Solar, and ECL datasets, we use both transformer encoders and MLPs for channel embedding. For ETTh1, ETTm1, ETTm2, and Weather, we use only MLP layers for channel embedding. For the Traffic dataset, we use only transformer encoders and adopt a channel-dependent approach to improve efficiency (though this is not necessary).
For datasets with fewer channels, such as the ETT series and Weather datasets, MLPs are sufficient for channel embedding. In contrast, for datasets with a larger number of channels, transformer encoders are preferable for enhancing model performance. Additionally, as the number of channels increases, a lower dropout rate should be set for better model stability.
Appendix B Detail information for long time series forecasting result
The detailed information from Table 2 is presented in Table 5:
Table 5: Detail result for AverageTime and LightAverageTime in Long-term time series forecasting.
Datasets Avg.Time LAvg.Time TimeXer CycleNet iTrans. RLinear PatchTST TimesNet
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.378 0.393 0.376 0.392 0.382 0.403 0.375 0.395 0.386 0.405 0.386 0.395 0.414 0.419 0.384 0.402
ETTh1 192 0.434 0.424 0.435 0.425 0.429 0.435 0.436 0.428 0.441 0.436 0.437 0.424 0.460 0.445 0.436 0.429
336 0.475 0.446 0.474 0.445 0.468 0.448 0.496 0.455 0.487 0.458 0.479 0.446 0.501 0.466 0.491 0.469
720 0.466 0.461 0.464 0.460 0.469 0.461 0.520 0.484 0.503 0.491 0.481 0.470 0.500 0.488 0.521 0.500
96 0.290 0.340 0.292 0.343 0.286 0.338 0.298 0.344 0.297 0.349 0.288 0.338 0.302 0.348 0.340 0.374
ETTh2 192 0.370 0.390 0.376 0.393 0.363 0.389 0.372 0.396 0.380 0.400 0.374 0.390 0.388 0.400 0.402 0.414
336 0.414 0.426 0.417 0.428 0.414 0.423 0.431 0.439 0.428 0.432 0.415 0.426 0.426 0.433 0.452 0.452
720 0.413 0.435 0.423 0.441 0.408 0.432 0.450 0.458 0.427 0.445 0.420 0.440 0.431 0.446 0.462 0.468
96 0.314 0.353 0.318 0.358 0.318 0.356 0.319 0.360 0.334 0.368 0.355 0.376 0.329 0.367 0.338 0.375
ETTm1 192 0.356 0.375 0.360 0.378 0.362 0.383 0.360 0.381 0.387 0.391 0.391 0.392 0.367 0.385 0.374 0.387
336 0.392 0.398 0.392 0.400 0.395 0.407 0.389 0.403 0.426 0.420 0.424 0.415 0.399 0.410 0.410 0.411
720 0.463 0.438 0.460 0.435 0.452 0.441 0.447 0.441 0.491 0.459 0.487 0.450 0.454 0.439 0.478 0.450
96 0.170 0.251 0.170 0.253 0.171 0.256 0.163 0.246 0.180 0.264 0.182 0.265 0.175 0.259 0.187 0.267
ETTm2 192 0.235 0.294 0.236 0.295 0.237 0.299 0.229 0.290 0.250 0.309 0.246 0.304 0.241 0.302 0.249 0.309
336 0.293 0.332 0.297 0.334 0.296 0.338 0.284 0.327 0.311 0.348 0.307 0.342 0.305 0.343 0.321 0.351
720 0.396 0.393 0.396 0.393 0.392 0.394 0.389 0.391 0.412 0.407 0.407 0.398 0.402 0.440 0.408 0.403
96 0.156 0.203 0.156 0.203 0.157 0.205 0.158 0.203 0.174 0.214 0.192 0.232 0.177 0.218 0.172 0.220
Weather 192 0.205 0.248 0.205 0.245 0.204 0.247 0.207 0.247 0.221 0.254 0.240 0.271 0.225 0.259 0.219 0.261
336 0.263 0.289 0.260 0.286 0.261 0.290 0.262 0.289 0.278 0.296 0.292 0.307 0.278 0.297 0.280 0.306
720 0.345 0.345 0.343 0.342 0.340 0.341 0.344 0.344 0.358 0.349 0.364 0.353 0.354 0.348 0.365 0.359
96 0.198 0.247 0.215 0.257 0.201 0.259 0.186 0.245 0.203 0.237 0.322 0.339 0.225 0.270 0.250 0.292
Solar 192 0.227 0.270 0.234 0.270 0.229 0.276 0.201 0.256 0.233 0.261 0.360 0.358 0.253 0.288 0.296 0.318
336 0.241 0.281 0.243 0.283 0.251 0.293 0.228 0.276 0.248 0.273 0.397 0.369 0.270 0.298 0.319 0.330
720 0.251 0.286 0.255 0.281 0.251 0.295 0.226 0.274 0.249 0.275 0.396 0.361 0.269 0.299 0.338 0.337
96 0.142 0.240 0.144 0.242 0.140 0.242 0.136 0.229 0.148 0.240 0.201 0.281 0.195 0.285 0.168 0.272
ECL 192 0.161 0.256 0.163 0.258 0.157 0.256 0.152 0.244 0.162 0.253 0.201 0.283 0.199 0.289 0.184 0.289
336 0.177 0.275 0.181 0.278 0.176 0.275 0.170 0.264 0.178 0.269 0.215 0.298 0.215 0.305 0.198 0.300
720 0.217 0.310 0.217 0.310 0.211 0.306 0.212 0.299 0.225 0.317 0.257 0.331 0.256 0.337 0.220 0.320
96 0.394 0.263 0.396 0.265 0.428 0.271 0.458 0.296 0.395 0.268 0.649 0.389 0.462 0.295 0.593 0.321
Traffic 192 0.415 0.274 0.418 0.276 0.448 0.282 0.457 0.294 0.417 0.276 0.601 0.366 0.466 0.296 0.617 0.336
336 0.428 0.280 0.430 0.281 0.473 0.289 0.470 0.299 0.433 0.283 0.609 0.369 0.482 0.304 0.629 0.336
720 0.457 0.297 0.460 0.300 0.516 0.307 0.502 0.314 0.467 0.302 0.647 0.387 0.514 0.322 0.640 0.350
Our model shows great robotness for all datasets, showing the effectivness of average framework for long-term time series forecasting.